# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 4.2 Learning Multi-Index Models with RCN
In this section, we demonstrate that our algorithmic approach becomes more efficient when the labeled examples are perturbed by uniform random noise. In particular, we study the performance of our approach in the RCN setting (see Definition 4.8). We show that Algorithm 3 can be modified into Algorithm 5, which achieves improved complexity in the RCN regime compared to the agnostic noise setting. The main result of this section is Theorem 4.9.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$. We say that $D$ is a RCN distribution for the class $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ if there exists a function $f\in\mathcal{F}$ and a row-stochastic matrix $\mathbf{H}\in\mathbb{R}^{\lvert\mathcal{Y}\rvert\times\lvert\mathcal{Y}\rvert}$ with $\mathbf{H}_{i,i}>\mathbf{H}_{i,j}+\gamma$ for all $i,j\in\mathcal{Y}$, for some $\gamma>0$, such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{x}),i}$. We refer to $\mathbf{H}$ as the confusion matrix of the distribution.
The key distinction that allows for this improvement is that, in the RCN setting, the output label depends solely on the ground truth label. Therefore, the moments of the noisy class regions still rely only on the subspace that characterizes the classifier. Consequently, any non-zero moment results in vectors that approximately lie within that subspace (see Proposition 4.10). The main result of this section is Theorem 4.9.
Input: Accuracy $\epsilon>0$, $\delta>0$ and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ that satisfies Definition 4.8.
Output: A hypothesis $h$ such that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
@@
      Draw a set $S$ of $N$ i.i.d. samples from $D$.
@@
      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
@@
    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as described in Definition 2.3, using $N$ i.i.d. samples from $D$, by finding the most frequent label at any set in $\mathcal{S}$.
@@
Theorem 4.9 (Learning with RCN).
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $\epsilon,\delta\in(0,1)$. Let $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$, where $\mathcal{F}$ is a class of $(m,{\zeta+\epsilon},\alpha,K,\Gamma,\tau,\sigma)$-well-behaved Multi-Index Models, as in Definition 3.2. If $\zeta\geq\mathrm{OPT}$ and $D$ satisfies Definition 4.8, then Algorithm 5 draws $N={(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}}$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(d,N,\lvert\mathcal{Y}\rvert)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
@@
We remark that for the case of multiclass linear classifiers, similarly to Proposition 4.2, using Proposition 2.4 we can prove that the class of multiclass linear classifiers is $(1,\zeta,\epsilon,K,O(\zeta)+O(\epsilon),\mathrm{poly}(\epsilon/K),O(K\sqrt{\log(K)}))$-well behaved. As a result, we obtain an algorithm for learning under RCN by applying Theorem 4.9. Moreover, note that in this case Algorithm 4 performs linear regression under Gaussian marginals (see Line 4). This task that can be efficiently reduced to mean estimation, specially to estimating $\operatorname*{\mathbf{E}}[\mathds{1}(y=i)]$ and $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$. Therefore, this step can be performed in polynomial time using $O(d)$ samples. Consequently, the final sample complexity of the algorithm scales with $d$ rather than $d^{O(1)}$, which establishes Theorem 1.8.
@@
Our proof of the Theorem 4.9 relies on the following proposition, which states that, given that our distribution is an RCN distribution with respect to a function in the class, all vectors returned by Algorithm 4 have an improved correlation with the ground truth subspace compared to the agnostic case (see Proposition 3.7).
Proposition 4.10.
@@
Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{3}/(\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2}))$, for $c$ a sufficiently small universal constant. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ and let $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ be a class of $(m,{\zeta+O(\epsilon)},\alpha,K,\Gamma,\tau,\sigma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ and let $W$ be a subspace such that $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$, with $K=\dim(W)$. Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according to $V$, as defined in Definitions 2.2 and 2.3.
@@
There exists $N=(dm)^{O(m)}{(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}}$ such that, if $D$ satisfies Definition 4.8 for $f$, $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ and $\zeta\geq\mathrm{OPT}$, then Algorithm 4 given $N$ i.i.d. samples from $D$, runs in $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time, and with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma/(\epsilon\sigma\alpha))$, such that: for any vector $\mathbf{v}\in\mathcal{E}$, there exists a unit vector $\mathbf{w}\in W$ for which
@@
Denote by $\mathbf{H}$ the stochastic matrix that perturbs the labels of $f$. Also let $\widehat{\mathbf{U}}$ be the matrix computed at Line 6 of Algorithm 5 and let $\eta^{2}=\mathrm{poly}(\epsilon\sigma\alpha/(mK\Gamma))$ be the $l_{2}^{2}$ error chosen in the polynomial-regression step, i.e., Line 4 of Algorithm 5.
We begin by showing that, for each subset of the partition, the computed vectors have only a small component in the subspace orthogonal to $W$. This, in turn, implies that any large vector must have a significant component in $W$.
@@
Fix $S\in\mathcal{S}$. If the number of samples that fall in $S$ is $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}/\delta\rvert)/\eta^{C}$ for a sufficiently large universal constant $C>0$, then for any $\mathbf{u}\in U_{S}$ and unit vector $\mathbf{v}\in W^{\perp}$ we have that $\lvert\mathbf{v}\cdot\mathbf{u}\rvert\leq\sqrt{2Km}\eta/\sigma$.
@@
Let $N_{S},S\in\mathcal{S}$, be the number of samples that fall in the set $S$. From the rotational invariance of the standard Gaussian, without loss of generality let $\mathbf{e}_{1},\dots,\mathbf{e}_{t}$ be a basis of $W^{\perp}$, $t\in\mathbb{Z}_{+}$. By B.2 and the union bound, we have that with $N_{S}=(dm)^{O(m)}\log(\lvert\mathcal{Y}\rvert/\delta)/{\eta^{O(1)}}$ i.i.d. samples and runtime $\mathrm{poly}(N_{S},d)$, we can find polynomials $p_{S,i}$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}]+\eta^{2}$. Hence, by averaging over $V$ conditioning on $S$, we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$, where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$.
Define $\hat{p}_{S,i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})],\hat{g}_{i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]$ for each $\beta\in\mathbb{N}^{d}$. Note that $\hat{p}_{S,i}(\beta)$ and $\hat{g}_{i}(\beta)$ are the Hermite coefficients of $p_{S,i}$ and $\mathds{1}(y=i)$, respectively. Since by the orthogonality of Hermite polynomials we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\in\mathbb{N}^{d}}(\hat{p}_{S,i}(\beta)-\hat{g}_{i}(\beta))^{2}$, it holds that $\sum_{\beta\in\mathbb{N}^{d},1\leq\|\beta\|_{1}\leq m}(\hat{p}_{S,i}(\beta)-\hat{g}_{i}(\beta))^{2}\leq\eta^{2}$.
Moreover, if $\beta_{i}\neq 0$, then for some $i\in[t]$
$\hat{g}_{i}(\beta)$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(y=i)\mid S]=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(f(\mathbf{x})=j)\mid S]$
@@
where we used the fact that the $\mathbf{x}\in S$ depends only on $\mathbf{x}^{V}$, and the fact that non-constant Hermite polynomials have mean zero under the Gaussian distribution. Therefore, we have that the sum of the Hermite coefficients $\hat{p}_{S,i}(\beta)$, for all $\beta\in\mathbb{N}^{d}$ such that $1\leq\|\beta\|_{1}\leq m$ and $\beta_{i}\neq 0$ for some $i\in[t]$, squared is less than $\eta^{2}$.
By fact A.2, for any $S\in\mathcal{S},i\in[K]$, we have that
$\mathbf{e}_{1}^{\top}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]\mathbf{e}_{1}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\nabla p_{S,i}(\mathbf{x})\cdot\mathbf{e}_{1})^{2}]$ $=\sum_{\beta\in\mathbb{N}^{d}}\beta_{1}(\hat{p}_{S,i}(\beta))^{2}$
@@
Hence, as $U_{S,i}$ has the eigenvectors of $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\nabla p_{S,i}(\mathbf{x}^{\perp V})\nabla p_{S,i}(\mathbf{x}^{\perp V})^{\top}\mid S]$ with eigenvalue greater than $\sigma^{2}/(2K)$, by Claim 4.12 we have the statement of the claim. ∎
@@
Let $\mathbf{M}\in\mathbb{R}^{d\times d}$ be a symmetric, psd matrix and let $\mathbf{v}\in\mathbb{R}^{d}$ be a vector such that $\mathbf{v}^{\top}\mathbf{M}\mathbf{v}\leq\epsilon$. Let $U$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalue greater than $\lambda$. Then, for every $\mathbf{u}\in U$, it holds that $\lvert\mathbf{u}\cdot\mathbf{v}\rvert\leq\sqrt{\epsilon/\lambda}$.

### v2

@@ 4.2 Learning Multi-Index Models with RCN
In this section, we demonstrate that our algorithmic approach becomes more efficient when the labeled examples are perturbed by uniform random noise. In particular, we study the performance of our approach in the RCN setting (see Definition 4.8). We show that Algorithm 3 can be modified into Algorithm 5, which achieves improved complexity in the RCN regime compared to the agnostic noise setting. The main result of this section is Theorem 4.9 (Theorem 1.11).
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$. We say that $D$ is an RCN distribution for the class $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ if there exists a function $f\in\mathcal{F}$ and a row-stochastic matrix $\mathbf{H}\in\mathbb{R}^{\lvert\mathcal{Y}\rvert\times\lvert\mathcal{Y}\rvert}$ with $\mathbf{H}_{i,i}>\mathbf{H}_{i,j}+\gamma$ for all $i,j\in\mathcal{Y}$, for some $\gamma>0$, such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{x}),i}$. We refer to $\mathbf{H}$ as the confusion matrix of the distribution.
The key distinction that enables us to have an improved algorithm (in terms of complexity) in the RCN setting is that the output label is determined solely by the ground truth label. Consequently, the moments of the observed labels (noisy labels) depend only on the subspace defining the classifier. This implies that any nonzero moment yields vectors that lie approximately within that subspace (see Proposition 4.10). The main result of this section is stated in Theorem 4.9.
Input: $\epsilon,\delta>0$ and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ that satisfies Definition 4.8 and whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
Output: A hypothesis $h$ such that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$, where $\mathrm{OPT}\coloneqq\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x}){\neq}y]$.
@@
      Draw a set $S_{t}$ of $N$ i.i.d. samples from $D$.
@@
      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S_{t}$).
@@
    Draw $N$ i.i.d. samples from $D$ and construct the piecewise constant classifier $h_{\mathcal{S}}$ as follows: For each $S\in\mathcal{S}$, assign a label that appears most frequently among the samples falling in $S$. Return $h_{\mathcal{S}}$.
@@
Theorem 4.9 (Learning MIMs with RCN).
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $\epsilon,\delta\in(0,1/2)$. Let $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$, where $\mathcal{F}$ is a class of $(m,\zeta+\epsilon,\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models, as in Definition 3.2. If $\zeta\geq\mathrm{OPT}$ and $D$ satisfies Definition 4.8, then Algorithm 5 draws $N=(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
@@
We remark that for the class of multiclass linear classifiers using Proposition 2.4, we can show that this class is $(1,\zeta,\epsilon,K,O(\zeta+\epsilon),\mathrm{poly}(\epsilon/K),O(K\sqrt{\log(K)}))$-well-behaved. As a result, by applying Theorem 4.9, we obtain an efficient algorithm for learning under Random Classification Noise (RCN). Moreover, for the class of multiclass linear classifiers Algorithm 4 performs linear regression under Gaussian marginals (see Line 4). This task that can be efficiently reduced to mean estimation, specially to estimating $\operatorname*{\mathbf{E}}[\mathds{1}(y=i)]$ and $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$. Since these mean estimates can be computed in polynomial time using $O(d)$ samples, the overall sample complexity of the algorithm scales as $d$, rather than $d^{O(1)}$. This establishes Theorem 1.11.
@@
Our proof of the Theorem 4.9 relies on the following proposition, which states that, given that our distribution is an RCN distribution with respect to a function in the class, any vector returned by Algorithm 4 has an improved correlation with the ground truth subspace compared to the agnostic case (see Proposition 3.7).
Proposition 4.10 (Structural Result for RCN).
@@
Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{3}/(\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2}))$, for $c$ a sufficiently small universal constant. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ and let $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ be a class of $(m,\zeta+O(\epsilon),\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Consider a function $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ and suppose that $f$ depends only on a $K$-dimensional subspace $W$, i.e., $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$. In addition, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$ with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$ of $V$ (as defined in Definitions 2.2 and 2.3).
@@
Then, there exists a sample size $N=(dm)^{O(m)}(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}$ such that if $D$ satisfies the RCN condition (see Definition 4.8) for $f$, and if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ and $\zeta\geq\mathrm{OPT}$, then Algorithm 4, when given $N$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(N)$ and, with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma/(\epsilon\sigma\alpha))$, such that for every vector $\mathbf{v}\in\mathcal{E}$, there exists a unit vector $\mathbf{w}\in W$ with
@@
Denote by $\mathbf{H}$ the stochastic matrix that perturbs the labels of $f$. Also let $\widehat{\mathbf{U}}$ be the matrix computed in Line 6 of Algorithm 5 and let $\eta^{2}=\mathrm{poly}(\epsilon\sigma\alpha/(mK\Gamma))$ be the $L_{2}^{2}$ error chosen in the polynomial-regression step, i.e., Line 4 of Algorithm 5.
We show that if the sample size is sufficiently large then for each region $S\in\mathcal{S}$ the associated set $U_{S}$ (see Line 5 of Algorithm 4) contains only vectors that have only a small component in the subspace orthogonal to $W$. This in turn implies that any large vector must have a significant component in $W$.
@@
Fix $S\in\mathcal{S}$. Suppose that the number of samples falling in $S$ satisfies $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}/\delta\rvert)/\eta^{C}$ for a sufficiently large universal constant $C>0$. Then, for any $\mathbf{u}\in U_{S}$ and any unit vector $\mathbf{v}\in W^{\perp}$ we have that $\lvert\mathbf{v}\cdot\mathbf{u}\rvert\leq\sqrt{2Km}\eta/\sigma$.
@@
Let $N_{S},S\in\mathcal{S}$, be the number of samples that land in the set $S$. From the rotational invariance of the standard Gaussian, without loss of generality let $\mathbf{e}_{1},\dots,\mathbf{e}_{t}$ be a basis of $W^{\perp}$, $t\in\mathbb{Z}_{+}$. By B.2 and the union bound, we have that with $N_{S}=(dm)^{O(m)}\log(\lvert\mathcal{Y}\rvert/\delta)/\eta^{O(1)}$ i.i.d. samples and runtime $\mathrm{poly}(N_{S},d)$, we can compute polynomials $p_{S,i}$ such that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}]+\eta^{2}\;.$
Hence, by averaging over $V$ conditioned on the cube $S$, we obtain that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}\;,$
where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$. Define, for each multi-index $\beta\in\mathbb{N}^{d}$, the Hermite coefficients of $p_{S,i}$ and $\mathds{1}(y=i)$ by $\widehat{p}_{S,i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})]$ and $\widehat{g}_{i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]$. By the orthogonality of Hermite polynomials, we have
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\in\mathbb{N}^{d}}(\widehat{p}_{S,i}(\beta)-\widehat{g}_{i}(\beta))^{2}\;.$
Thus, restricting the sum to multi-indices $\beta$ with $1\leq\|\beta\|_{1}\leq m$, it follows that $\sum_{\beta\in\mathbb{N}^{d},1\leq\|\beta\|_{1}\leq m}(\widehat{p}_{S,i}(\beta)-\widehat{g}_{i}(\beta))^{2}\leq\eta^{2}$. Moreover, if $\beta_{i}\neq 0$, then for some $i\in[t]$, it holds
$\widehat{g}_{i}(\beta)$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(y=i)\mid S]=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(f(\mathbf{x})=j)\mid S]$
@@
where we used the fact that each $\mathbf{x}\in S$ depends only on $\mathbf{x}^{V}$, and any non-constant Hermite polynomial has zero mean under the standard Gaussian distribution. Thus, let $E_{error}$ denote the set of multi-indices $\beta$ for which the corresponding Hermite coefficient of $\mathds{1}(y=i)$ is zero, i.e., $\widehat{g}_{i}(\beta)=0$. Then, we have that the sum of the Hermite coefficients, restricted to multi-indexes that in set $E_{error}$ is bounded by
$\sum_{\beta\in E_{error}}\widehat{p}_{S,i}(\beta)^{2}\leq\eta^{2}\;.$
By A.2, for any $S\in\mathcal{S},i\in\mathcal{Y}$, we have that
$\mathbf{e}_{1}^{\top}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]\mathbf{e}_{1}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\nabla p_{S,i}(\mathbf{x})\cdot\mathbf{e}_{1})^{2}]$ $=\sum_{\beta\in\mathbb{N}^{d}}\beta_{1}(\widehat{p}_{S,i}(\beta))^{2}$
@@
Hence, since $U_{S,i}$ contains the eigenvectors of $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]$ with eigenvalue greater than $\sigma^{2}/(2K)$, using Claim 4.12 we complete the proof. ∎
@@
Let $\mathbf{M}\in\mathbb{R}^{d\times d}$ be a symmetric, PSD matrix and let $\mathbf{v}\in\mathbb{R}^{d}$ be a unit vector such that $\mathbf{v}^{\top}\mathbf{M}\mathbf{v}\leq\epsilon$. Let $U$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalue at least $\lambda$. Then, for every $\mathbf{u}\in U$, it holds that $\lvert\mathbf{u}\cdot\mathbf{v}\rvert\leq\sqrt{\epsilon/\lambda}$.